Adult tickets for a play cost $15 and child tickets cost $13. If there were 26 people at a performance and the theater collected $364 from ticket sales, how many adults and how many children attended the play?
VERY bad with word problems

Answer by checkley75(3666)

15A+13C=364
A+C=26 OR A=26-C
NOW REPLACE A WITH(26-C) IN THE FIRST EQUATION.
15(26-C)+13C=364
390-15C+13C=364
-2C=364-390
-2C=-26
C=-26/-2
C=13 CHILDREN TICKETS SOLD.
26-13=13 ADULT TICKETS SOLD.
PROOF
15*13+13*13=364
195+169=364
364=364